Are the proportions of road accidents similar in various highways of Bangladesh?                    

Highways

1

2

3

4

Total

No. of road accidents (Oi)

50

42

32

82

206



1
Expert's answer
2021-11-26T05:13:45-0500

The hypothesis tested in this question are,

"H_0:"  Proportions of road accidents are similar.

"Against"  

"H_1:"  Proportions of road accidents are not similar.

To perform this test we shall the chi-square goodness of fit test.

If the null hypothesis is true, then the proportion for each highway should be the same. Since there are 4 highways, the proportion for each highway is "p=1\/4" .

We proceed to find the expected count for each highway as follows,

"E_k=n*p" for "k=1,2,3,4" and "n=206"

"E_1=206*1\/4=51.5"

"E_2=206*1\/4=51.5"

"E_3=206*1\/4=51.5"

"E_5=206*1\/4=51.5"

The test statistic is given as,

"\chi^2_c=\displaystyle \sum^4_{k=1}(O_k-E_k)^2/E_k"

"\chi^2_c=(50-51.5)^2/51.5+(42-51.5)^2/51.5+(32-51.5)^2/51.5+(82-51.5)^2/51.5= 27.24272"

"\chi^2_c" is compared with the chi square table value at "\alpha=5\%" with "(k-1)=(4-1)=3", where "k" is the number of highways.

The table value is given as, "\chi^2_{\alpha, 3}=\chi^2_{0.05,3}=7.81473" and the null hypothesis is rejected if "\chi^2_{c}\gt \chi^2_{0.05,3}".

Since "\chi^2_{c}=27.24272\gt \chi^2_{0.05.3}=7.81473" the null hypothesis is rejected and we conclude that there is not enough evidence to show that the proportions of road accidents are similar in various highways of Bangladesh at 5% level of significance.